A company manufactures batteries in batches of 8 and there is a 3% rate of defects. Find the mean number of defects per batch.
Answer #1

Solution:

p = 3% = 0.03

n = 8

X be the number of defects in this batch of 8

X follows Binomial(n = 8 , p = 0.03)

Mean of the binomial distribution is

Mean = n * p = 8 * 0.03 = 0.24

Answer : 0.24